Systems Staff Meeting Notes 7/31/00

Systems Staff Meeting Notes
July 31, 2000

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Debra Harris, Chris Leopold, Bobbie Meng, Suzy Shaw, Geoffrey Stewart, Ying Tang
Announcements
  1. Chris Leopold is joining the meeting today. He works at IFAS, but will be the off-site backup for Bill while is on leave of absence.
  2. First Data Czar Act: Web Control - creation of myflorida.com, a single point of contact on the Web for all of the state resources. UF is trying to coordinate with the state. A network of UF Web managers is being developed in an attempt to mesh with the state standards. The privacy policy will be the first issue on the table for the managers.
Review Minutes of 7/24/00
  1. #4 Software - Dialog Interface - Winston will email John Ashcraft, not Freund, regarding status.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Visors - the order for the other 15 has been impossible. We've given up on Best Buy, as there is no telling what is going on with the order. From now on we'll be buying them directly from Handspring. This will result in it taking longer to receive them. Five ice-colored ones were recently received, apparently the tail end of an old order. Bill would like to know immediately if we receive any boxes from Best Buy.
  2. Geoffrey found out that Handspring will do repair work, but will not process a repair on the Visors without a credit card number. Betty Mitchell has a VISA card; she will be the contact.
  3. Security cameras have been ordered for DLC.
  4. The HSS laptop that Will had returned for repair to Gateway came back with less memory than when it left. We sent it back for them to crank up the memory.
Liaisons
  1. Meeting scheduled for this Thursday.
  2. Scheduling appointments for repair with liaisons - Geoffrey had success with Melody in DLC. She watched the whole process, diligently taking notes. However, he had little interest on the 3rd floor.
Networking
  1. AC outage this weekend - North tower was down, but all is back up again as of this morning.
  2. Terabyte server - still in progress.
  3. Closet micros - ready to go in MSL, East & West.
  4. East 3rd floor rewire for RS/DLC - people have been re-podded. Two new pods have appeared on the 3rd floor. Laydene is currently the only inhabitant on the 2nd floor. The phones have not been re-situated. All of the 3rd floor is networked and the AC is working. IP addresses moved with machines, so no changes were necessary. DLC move won't happen until September due to some problem with the carpet. Tim & Judy from Digital plan on bringing a crew over some time this week. Bobbie thought they should hold off, since things aren't actually ready for them. Bill wants them to interact directly with Facilities (not us) since Systems had nothing to do with the scheduling.
  5. Branch wiring adjustments - on hold until East project is complete.
  6. Wireless - awaiting Facilities; no funding for conduit.
Software Projects Status
  1. Systems wiring - Suzy installed VISIO and did a preliminary diagram of West. She needs to get access to the IR database. Ying can install it on her machine.
  2. RS - still in progress.
  3. DLC - tracking database still in progress. Erich was out of town last week, but he emailed Ying with two new requests.
DP Coordinator
  1. Students will start to disappear in two weeks - summer term ends on August 11.
Calendar
  1. Winston & Hilton are both out today.
  2. Ying might be out Aug. 14.
  3. Cynthia will be out Aug. 4 & Aug. 11.


Back to the Systems Homepage

Last updated August 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu